# Supercooled phase transitions in conformal dark sectors explain NANOGrav data

### v2

@@ Abstract
Consider imposing a classical conformal symmetry in both the dark and visible sectors, i.e., $\mu^{2}_{h}=0$, and requiring a large hierarchy between the vevs, $v_{\sigma}\gg v$, associated with a large hierarchy in the scalar mass spectrum, $M_{h_{2}}\gg M_{h_{1}}$ Gonçalves et al. (2025). (Throughout, we fix the mass of the SM-like Higgs boson to its experimentally measured value $M_{h_{1}}=125.11~{}\mathrm{GeV}$ Aad et al. (2023)) In this case, the scalon emerges from the heavy singlet $\sigma$ field. This is in variance with a multi-Higgs-doublet scenario which necessarily implies that the scalon is the lightest particle in the scalar spectrum Gildener and Weinberg (1976). Note that a large mass hierarchy between the scalon and the SM Higgs is protected against large radiative corrections by means of a small portal coupling $|\lambda_{\sigma h}|\sim(v/v_{\sigma})^{2}$. In this scenario, GW signals with peak frequencies above a mHz are obtained Gonçalves et al. (2025).
@@
On the other hand, an FOPT induced by the breaking of the gauged $\mathrm{U(1)}^{\prime}$ symmetry at a MeV scale could produce a GW signal in the nHz range relevant to PTA observations. However, in this case, with the opposite hierarchy $v\gg v_{\sigma}$, corresponding to $M_{h_{1}}\gg M_{h_{2}}$, the scalon emerges from $\mathcal{H}$. Now there are two possibilities: 1) If $\lambda_{h}<y_{t}$, where $y_{t}$ is the top Yukawa coupling, the top quark loop is the dominant contribution to the Higgs mass, which comes with a negative sign and alters its vacuum stability, such that no valid solution exists. 2) If $\lambda_{h}>y_{t}$, the scalar loop is the dominant contribution and valid solutions are possible. However, as shown in Ref. Elias et al. (2003), the Higgs mass predicted by minimizing the effective potential exceeds 200 GeV, which is inconsistent with experiment.
@@
We avoid this problem by resorting to a novel approach where only the dark sector is conformal i.e., the $\sigma^{\dagger}\sigma$ term is absent, but not the visible one ($\mu^{2}_{h}\not=0$). This is necessary to ensure proper EW symmetry breaking with $M_{h_{1}}\gg M_{h_{2}}$. The Coleman-Weinberg (CW) mechanism Coleman and Weinberg (1973) of radiative $\mathrm{U(1)}^{\prime}$ symmetry breaking is analyzed at one-loop order. Following Ref. Gonçalves et al. (2025), the potential in Eq. (1) is supplemented by the one-loop CW potential Coleman and Weinberg (1973), i.e., $V=V_{0}+V_{\mathrm{CW}}$, enabling us to also compute the mass spectrum at one loop. This procedure fixes the values of $\lambda_{h}$, $\lambda_{\sigma}$, $\mu^{2}_{h}$ and the singlet vev $v_{\sigma}$, leaving the scalon mass $M_{h_{2}}$ and the scalar mixing $\lambda_{\sigma h}$ as the physical input parameters of the scalar sector. In the gauge sector, the additional Abelian gauge field mixes kinetically with the hypercharge $\mathrm{U(1)}_{\rm Y}$ field. This can be parameterized by a gauge coupling $g_{12}$ in addition to the $\mathrm{U(1)}^{\prime}$ gauge coupling $g_{L}$. We fix the value of $g_{12}$ as described below. Hence, the only additional free parameter arising from the gauge sector is $g_{L}$. In the renormalization group (RG) evolution, we require that $\lambda_{h}$ remains positive and that all couplings are perturbative up to a renormalization scale $\mu=10^{8}~{}\mathrm{GeV}$, which sets the scale to which our model is valid.
@@
Since we focus on phase transition temperatures much lower than the EW scale, we do not consider contributions from $W^{\pm}$, $\mathrm{Z^{0}}$ and the top quark in the RG evolution. Also, we only consider the $\sigma$ direction to be relevant for the FOPT as it decouples from the Higgs direction due to the strong scale hierarchy $M_{h_{1}}\gg M_{h_{2}}$. The field-dependent RG scale is chosen to be $\mu=\mathrm{max}[M_{\mathrm{Z^{\prime}}}(\phi_{\sigma}),\pi T]$, where $M_{\mathrm{Z^{\prime}}}(\phi_{\sigma})$ is the $\mathrm{Z^{\prime}}$ field-dependent mass,
@@
FOPTs in conformal dark sectors. FOPTs in conformal models are typically supercooled as the potential barrier between the true and false vacuum persists as $T\rightarrow 0$ allowing for percolation to occur at arbitrarily low temperatures.
Our scenario overcomes the issues highlighted in Ref. Athron et al. (2024). First, percolation is always possible. In non-conformal models, such as the type studied in Ref. Athron et al. (2024), the action as a function of temperature exhibits a $U$-shaped behavior (see e.g., Fig. 4 of Ref. Athron et al. (2023)), which bounds the nucleation temperature from below. In contrast, conformal models have an action that asymptotically approaches zero as $T\rightarrow 0$ Levi et al. (2023). Second, we find that for $g_{L}$ of ${\cal{O}}(0.1)$ supercooled FOPTs complete Gonçalves et al. (2025). Third, the conformal nature of the potential prevents $T_{\mathrm{RH}}\approx M$, where $M$ is the mass scale of dark sector. Since the potential barrier persists as $T\rightarrow 0$, the FOPT can be delayed to temperatures well below $T_{c}$, which defines the energy scale driving the transition. Consequently, as long as extreme supercooling is avoided ($\alpha\lesssim 10^{8}$), $T_{\mathrm{RH}}$ deviates from $T_{p}$ by at most a factor of $\mathcal{O}(1-100)$, ensuring that $T_{\mathrm{RH}}<M$ (as also noted in Ref. Ellis et al. (2020), where $T_{\mathrm{RH}}\sim 0.1M_{\mathrm{Z^{\prime}}}$ for a conformal $\mathrm{U}(1)_{\mathrm{B-L}}$ model). In this regime, Daisy corrections are negligible.
@@
We perform a scan over the model parameters at $\mu=M_{Z^{0}}$ with a focus on the parameter space capable of explaining the NANOGrav data. Since we focus on MeV scale FOPTs, we sample $M_{h_{2}}$ logarithmically in the range $M_{h_{2}}=[10^{-4},10]~{}\mathrm{GeV}$. The scalar mixing is scanned logarithmically in the range $\lambda_{\sigma h}=[10^{-15},10^{-10}]$, with the upper bound chosen so that $v_{\sigma}$ is small enough to yield a signal at PTAs. The gauge coupling is sampled linearly in the range $g_{L}=[0.25,1.5]$. Figure 1 displays scatter plots of the SGWB peak amplitude versus peak frequency for $g_{L}$ (top panel), $\lambda_{\sigma}$ (middle panel) and $M_{\mathrm{Z^{\prime}}}$ (bottom panel). Also shown is the NANOGrav signal Agazie et al. (2023) and the sensitivity of the experiments, GAIA, SKA, and THEIA Garcia-Bellido et al. (2021); Weltman et al. (2020). Our analysis indicates that the peak frequency is primarily determined by $M_{\mathrm{Z^{\prime}}}$, with smaller masses corresponding to lower frequencies. $M_{h_{2}}$ exhibits a similar correlation and is not shown. The couplings $\lambda_{\sigma}$ and $g_{L}$ influence the peak amplitude, with smaller values resulting in higher amplitudes. This is because they affect the location of the minimum of the effective potential and its depth Gonçalves et al. (2025). We also require $T_{\mathrm{RH}}<M$ to prevent extreme supercooling, as greater supercooling results in higher $T_{\mathrm{RH}}$, which shifts the GW spectrum to higher frequencies. The isolated points with $h^{2}\Omega_{\mathrm{GW}}^{\mathrm{peak}}<10^{-11}$ correspond to non-supercooled FOPTs with $\alpha<0.1$ and $g_{L}>1.2$.
@@
$g_{L}$ $\lambda_{\sigma h}/10^{-11}$ $M_{h_{2}}$ $M_{\mathrm{Z^{\prime}}}$ $T_{\mathrm{RH}}$ $\alpha/10^{5}$ $\beta/H$ $\gamma_{\mathrm{BHB}}$ $A_{\mathrm{BHB}}$
$0.60$ $4.50$ $12.4$ $107.3$ $11.7$ $2.60$ $39.5$ $4.50$ $10^{-15.4}$
Table 1: Parameters for the best fit point (including the contribution from SMBHBs) shown in Fig. 2. $M_{h_{2}}$, $M_{\mathrm{Z^{\prime}}}$ and $T_{\mathrm{RH}}$ are in units of MeV.
To determine the 68% and 95% confidence level (CL) parameter space regions allowed by the NANOGrav 15-year dataset, we employ the PTArcade package in its default ceffyl configuration Mitridate et al. (2023); Lamb et al. (2023). Figure 3 shows allowed regions assuming either that the model fully explains the signal (labeled "PT") or that the contribution from SMBHBs is included (labeled "PT+SMBHBs"). For the SMBHB signal, we adopt the default settings of PTArcade, choosing an uniform prior $\gamma_{\mathrm{BHB}}=[0,7]$ and a log-uniform prior $\mathrm{log_{10}}A_{\mathrm{BHB}}=[-18,-14]$. NANOGrav data favor $\mathrm{Z^{\prime}}$ and $h_{2}$ masses in the 10–100 MeV range at 68% CL, with larger masses disfavored due to the resultant higher peak frequencies. We observe a strong correlation between $\lambda_{\sigma}$ and $g_{L}$. Smaller couplings are favored because they yield larger peak amplitudes. Without the SMBHB contribution, we find $g_{L}\in[0.60,0.89]$ and $\lambda_{\sigma}\in[0.04,0.21]$ at 68% CL. As the top-right panel shows, the scalar mixing $\lambda_{\sigma h}$ is uncorrelated with the SGWB signal. As expected, including the SMBHB signal widens the allowed parameter regions, although in most cases the widening is minimal because the corresponding parameters show weak correlations with either the amplitude, frequency, or both.
@@
Summary. We showed that a dark conformal extension of the SM augmented by a dark $\mathrm{U(1)}^{\prime}$ symmetry can produce a supercooled FOPT that explains the GW signal observed by NANOGrav. The model is successful provided the FOPT is not extremely supercooled ($\alpha\lesssim 10^{8}$). The pitfalls noted in Ref. Athron et al. (2024) are overcome by ensuring that the phase transition completes and that the dark sector remains non-relativistic after reheating. We also obtained model parameter values favored by the NANOGrav 15-year data. Specifically, dark sector masses of $\mathcal{O}(1-100)~{}\mathrm{MeV}$, couplings $g_{L}$ and $\lambda_{\sigma}$ of $\mathcal{O}(0.1)$ are favored.
@@ Appendix A Appendix: Renormalization group equations
$\begin{split}&\beta^{(1)}(g_{L})=\frac{41}{10}g_{12}^{2}g_{L}+\frac{g_{L}^{3}}{3}\,,\end{split}$ (5)
$\begin{split}&\beta^{(1)}(g_{1})=\frac{41}{10}g_{1}(g_{1}^{2}+g^{2}_{12})\,,\end{split}$ (6)
$\begin{split}&\beta^{(1)}(g_{2})=-\frac{19}{6}g^{2}_{2}\,,\end{split}$ (7)
$\begin{split}&\beta^{(1)}(g_{3})=-7g^{2}_{3}\,,\end{split}$ (8)
$\begin{split}&\beta^{(1)}(g_{12})=\frac{41}{10}g_{12}(g_{1}^{2}+g_{12}^{2})+\frac{1}{2}g_{12}g_{L}^{2}\,,\end{split}$ (9)
$\begin{split}&\beta^{(1)}(y_{t})=\frac{3}{2}y_{t}^{3}+y_{t}\Big{[}-\frac{17}{20}g_{1}^{2}-\frac{17}{20}g_{12}^{2}-\frac{9}{4}g_{2}^{2}-\\ &\hskip 50.00008pt8g_{3}^{2}+3yt^{2}\Big{]}\,,\end{split}$ (10)
$\begin{split}&\beta^{(1)}(\lambda_{h})=\frac{27}{200}g_{1}^{4}+\frac{9}{20}g_{1}^{2}g_{2}^{2}+\frac{9}{8}g_{2}^{4}-\frac{9}{5}g_{1}^{2}\lambda_{h}-\\ &\hskip 50.00008pt\frac{9}{5}g_{12}^{2}\lambda_{h}-9g_{2}^{2}\lambda_{h}+24\lambda_{h}^{2}+\lambda_{\sigma h}^{2}+\\ &\hskip 50.00008pt12\lambda_{h}y_{t}^{2}-6y_{t}^{4}\,,\end{split}$ (11)
$\begin{split}&\beta^{(1)}(\lambda_{\sigma})=2\Big{[}3g_{L}^{4}-6g_{L}^{2}\lambda_{\sigma}+10\lambda_{\sigma}^{2}+\lambda_{\sigma h}^{2}\Big{]}\,,\end{split}$ (12)
$\begin{split}&\beta^{(1)}(\lambda_{\sigma h})=\frac{\lambda_{\sigma h}}{10}\Big{[}-9g_{1}^{2}-9g_{12}^{2}-45g_{2}^{2}-60g_{L}^{2}+\\ &\hskip 50.00008pt120\lambda_{h}+80\lambda_{\sigma}+40\lambda_{\sigma h}+60y_{t}^{2}\Big{]}\,,\end{split}$ (13)

### v3

@@ Abstract
Consider imposing a classical conformal symmetry in both the dark and visible sectors, i.e., $\mu^{2}_{h}=0$, and requiring a large hierarchy between the vevs, $v_{\sigma}\gg v$, associated with a large hierarchy in the scalar mass spectrum, $M_{h_{2}}\gg M_{h_{1}}$ Gonçalves et al. (2025). (Throughout, we fix the mass of the SM-like Higgs boson to its experimentally measured value $M_{h_{1}}=125.11~\mathrm{GeV}$ Aad et al. (2023).) In this case, the scalon emerges from the heavy singlet $\sigma$ field. This is in variance with a multi-Higgs-doublet scenario which necessarily implies that the scalon is the lightest particle in the scalar spectrum Gildener and Weinberg (1976). Note that a large mass hierarchy between the scalon and the SM Higgs is protected against large radiative corrections by means of a small portal coupling $|\lambda_{\sigma h}|\sim(v/v_{\sigma})^{2}$. In this scenario, GW signals with peak frequencies above a mHz are obtained Gonçalves et al. (2025).
@@
On the other hand, an FOPT induced by the breaking of the gauged $\mathrm{U(1)}^{\prime}$ symmetry at an MeV scale could produce a GW signal in the nHz range relevant to PTA observations. However, in this case, with the opposite hierarchy $v\gg v_{\sigma}$, corresponding to $M_{h_{1}}\gg M_{h_{2}}$, the scalon emerges from $\mathcal{H}$. Now there are two possibilities: 1) If $\lambda_{h}<y_{t}$, where $y_{t}$ is the top Yukawa coupling, the top quark loop is the dominant contribution to the Higgs mass, which comes with a negative sign and alters its vacuum stability, such that no valid solution exists. 2) If $\lambda_{h}>y_{t}$, the scalar loop is the dominant contribution and valid solutions are possible. However, as shown in Ref. Elias et al. (2003), the Higgs mass predicted by minimizing the effective potential exceeds 200 GeV, which is inconsistent with experiment.
@@
We avoid this problem by resorting to a novel approach where only the dark sector is conformal i.e., the $\sigma^{\dagger}\sigma$ term is absent, but not the visible one ($\mu^{2}_{h}\not=0$). This is necessary to ensure proper EW symmetry breaking with $M_{h_{1}}\gg M_{h_{2}}$. The Coleman-Weinberg (CW) mechanism Coleman and Weinberg (1973) of radiative $\mathrm{U(1)}^{\prime}$ symmetry breaking is analyzed at one-loop order. Following Ref. Gonçalves et al. (2025), the potential in Eq. (1) is supplemented by the one-loop CW potential Coleman and Weinberg (1973), i.e., $V=V_{0}+V_{\mathrm{CW}}$, enabling us to also compute the mass spectrum at one loop. This procedure fixes the values of $\lambda_{h}$, $\lambda_{\sigma}$, $\mu^{2}_{h}$ and the singlet vev $v_{\sigma}$, leaving the scalon mass $M_{h_{2}}$ and the scalar mixing $\lambda_{\sigma h}$ as the physical input parameters of the scalar sector. However, the role of $\lambda_{\sigma h}$ is rendered insignificant by the quadratic Higgs term. In the gauge sector, the additional Abelian gauge field mixes kinetically with the hypercharge $\mathrm{U(1)}_{\rm Y}$ field. This can be parameterized by a gauge coupling $g_{12}$ in addition to the $\mathrm{U(1)}^{\prime}$ gauge coupling $g_{L}$. We fix the value of $g_{12}$ as described below. Hence, the only additional free parameter arising from the gauge sector is $g_{L}$. In the renormalization group (RG) evolution, we require that $\lambda_{h}$ remains positive and that all couplings are perturbative up to a renormalization scale $\mu=10^{8}~\mathrm{GeV}$, which sets the scale to which our model is valid.
@@
Since we focus on phase transition temperatures much lower than the EW scale, we do not consider contributions from $W^{\pm}$ and $\mathrm{Z^{0}}$ in the RG evolution. Also, we only consider the $\sigma$ direction to be relevant for the FOPT as it decouples from the Higgs direction due to the strong scale hierarchy $M_{h_{1}}\gg M_{h_{2}}$. The field-dependent RG scale is chosen to be $\mu=\mathrm{max}[M_{\mathrm{Z^{\prime}}}(\phi_{\sigma}),\pi T]$, where $M_{\mathrm{Z^{\prime}}}(\phi_{\sigma})$ is the field-dependent $\mathrm{Z^{\prime}}$ mass,
@@
FOPTs in conformal dark sectors.
Figure 1: Ratio of the Euclidean action to temperature $S_{3}/T$ as a function of $T$ for the best-fit point in Table 1. The vertical dotted lines indicate the critical temperature $T_{c}$, nucleation temperature $T_{n}$, percolation temperature $T_{p}$, and completion temperature $T_{f}$.
Figure 2: Scatter plot of the reheating temperature $T_{\mathrm{RH}}$ as a function of the $\mathrm{Z^{\prime}}$ mass.
FOPTs in conformal models are typically supercooled as the potential barrier between the true and false vacuum persists as $T\rightarrow 0$ allowing for percolation to occur at arbitrarily low temperatures.
Our scenario overcomes the issues highlighted in Ref. Athron et al. (2024). First, in conformal models percolation is always possible. We have explicitly verified that the volume of the false vacuum is decreasing at the percolation temperature $T_{p}$ by requiring
$H(T)\left(3+T\frac{dI}{dT}\right)\Biggr{|}_{\begin{subarray}{c}T=T_{p}\end{subarray}}<0\,$, (4)
where $H(T)$ is the Hubble parameter and $I(T)$ is the true vacuum volume per unit comoving volume. As discussed in Ref. Gonçalves et al. (2025), this condition may not necessarily be satisfied at $T_{p}$, but at some lower temperature. However, here we only consider FOPTs for which the percolation condition is fulfilled at $T=T_{p}$. Second, in non-conformal models, such as the type studied in Ref. Athron et al. (2024), the three-dimensional Euclidean action as a function of temperature exhibits a $U$-shaped behavior (see e.g., Fig. 3 of Ref. Levi et al. (2023)), which bounds the percolation temperature from below. In contrast, conformal models have an action that asymptotically approaches zero as $T\rightarrow 0$; see Fig. 1. Third, we have checked that the potential remains bounded from below at the FOPT completion temperature $T_{f}$, at which the true vacuum occupies 99% of the Universe. Fourth, the conformal nature of the potential prevents $T_{\mathrm{RH}}\approx M$, where $M$ is the mass scale of dark sector. Since the potential barrier persists as $T\rightarrow 0$, the FOPT can be delayed to temperatures well below $T_{c}$, which defines the energy scale driving the transition. Consequently, as long as extreme supercooling is avoided ($\alpha\lesssim 10^{8}$), $T_{\mathrm{RH}}$ deviates from $T_{p}$ by at most a factor of $\mathcal{O}(1-100)$, ensuring that $T_{\mathrm{RH}}<M$; see Fig. 2. (It was also noted in Ref. Ellis et al. (2020) that $T_{\mathrm{RH}}\sim 0.1M_{\mathrm{Z^{\prime}}}$ for a conformal $\mathrm{U}(1)_{\mathrm{B-L}}$ model.) In this regime, Daisy corrections are negligible.
@@
We perform a scan over the model parameters at $\mu=M_{\mathrm{Z^{0}}}$ with a focus on the parameter space capable of explaining the NANOGrav data. Since we focus on MeV scale FOPTs, we sample $M_{h_{2}}$ logarithmically in the range $M_{h_{2}}=[0.1,100]~\mathrm{MeV}$. We require $|\lambda_{\sigma h}|<10^{-10}$ so that $v_{\sigma}$ is small enough to yield a signal at PTAs. For these values of $\lambda_{\sigma h}$, the impact on the FOPT is negligible. The gauge coupling is sampled linearly in the range $g_{L}=[0.25,1.5]$. In what follows, all parameters are evaluated at $\mu=0.1~\mathrm{MeV}$ after RG evolution. Figure 3 displays scatter plots of the SGWB peak amplitude versus peak frequency for $g_{L}$ (top panel), $\lambda_{\sigma}$ (middle panel) and $M_{\mathrm{Z^{\prime}}}$ (bottom panel). Also shown is the NANOGrav signal Agazie et al. (2023) and the sensitivity of the experiments, GAIA, SKA, and THEIA Garcia-Bellido et al. (2021); Weltman et al. (2020). Our analysis indicates that the peak frequency is primarily determined by $M_{\mathrm{Z^{\prime}}}$, with smaller masses corresponding to lower frequencies. $M_{h_{2}}$ exhibits a similar correlation and is not shown. The couplings $\lambda_{\sigma}$ and $g_{L}$ influence the peak amplitude, with smaller values resulting in higher amplitudes. This is because they affect the location of the minimum of the effective potential and its depth Gonçalves et al. (2025). We also require $T_{\mathrm{RH}}<M$ to prevent extreme supercooling, as greater supercooling results in higher $T_{\mathrm{RH}}$, which shifts the GW spectrum to higher frequencies. The isolated points with $h^{2}\Omega_{\mathrm{GW}}^{\mathrm{peak}}<10^{-11}$ correspond to non-supercooled FOPTs with $\alpha<0.1$ and $g_{L}>1.1$.
@@
$g_{L}$ $M_{h_{2}}$ $M_{\mathrm{Z^{\prime}}}$ $T_{\mathrm{RH}}$ $\alpha/10^{5}$ $\beta/H$ $\gamma_{\mathrm{BHB}}$ $A_{\mathrm{BHB}}$
$0.59$ $12.4$ $107.3$ $11.7$ $2.60$ $39.5$ $4.50$ $10^{-15.4}$
Table 1: Parameters for the best fit point (including the contribution from SMBHBs) shown in Fig. 4. $M_{h_{2}}$, $M_{\mathrm{Z^{\prime}}}$ and $T_{\mathrm{RH}}$ are in units of MeV, and evaluated at $\mu=0.1$ MeV.
To determine the 68% and 95% confidence level (CL) parameter space regions allowed by the NANOGrav 15-year dataset, we employ the PTArcade package in its default ceffyl configuration Mitridate et al. (2023); Lamb et al. (2023). Figure 5 shows allowed regions assuming either that the model fully explains the signal (labeled "PT") or that the contribution from SMBHBs is included (labeled "PT+SMBHBs"). For the SMBHB signal, we adopt the default settings of PTArcade, choosing a uniform prior $\gamma_{\mathrm{BHB}}=[0,7]$ and a log-uniform prior $\mathrm{log_{10}}A_{\mathrm{BHB}}=[-18,-14]$. NANOGrav data favor $\mathrm{Z^{\prime}}$ and $h_{2}$ masses in the 10–100 MeV range at 68% CL, with larger masses disfavored due to the resultant higher peak frequencies. We observe a strong correlation between $\lambda_{\sigma}$ and $g_{L}$. Smaller magnitudes are favored because they yield larger peak amplitudes. Without the SMBHB contribution, we find $g_{L}\in[0.57,0.86]$ and $\lambda_{\sigma}\in[-0.062,-0.028]$ at 68% CL. As expected, including the SMBHB signal widens the allowed parameter regions, although in most cases the widening is minimal because the corresponding parameters show weak correlations with either the amplitude, frequency, or both.
@@
Summary. We showed that a dark conformal extension of the SM augmented by a dark $\mathrm{U(1)}^{\prime}$ symmetry can produce a supercooled FOPT that explains the GW signal observed by NANOGrav. The model is successful provided the FOPT is not extremely supercooled ($\alpha\lesssim 10^{8}$). The pitfalls noted in Ref. Athron et al. (2024) are overcome by ensuring that the phase transition completes and that the dark sector remains non-relativistic after reheating. We also obtained model parameter values favored by the NANOGrav 15-year data. Specifically, dark sector masses of $\mathcal{O}(1-100)~\mathrm{MeV}$, couplings $g_{L}$ of $\mathcal{O}(0.1)$ and $\lambda_{\sigma}$ of $\mathcal{O}(0.01)$ are favored. Interestingly, NANOGrav data require $|\lambda_{\sigma h}|<10^{-10}$.
@@ Appendix A Appendix: Renormalization group equations
[…]